OpenStudy (anonymous):

help?!?!!??(medal will be given) Determine if triangle DEF with coordinates D (2, 1), E (3, 5), and F (6, 2) is an equilateral triangle. Use evidence to support your claim. If it is not an equilateral triangle, what changes can be made to make it equilateral? Be specific.

OpenStudy (anonymous):

@Directrix help plz

OpenStudy (anonymous):

@myininaya help plz

OpenStudy (anonymous):

if its equilateral triangle all angles are 60deg no coordinates required for it.

OpenStudy (anonymous):

OR if all sides are equal , use distance formula to see if all sides r equal

Directrix (directrix):

Use the distance formula to get the lengths of segments DE, EF, and FD. The formula is attached. Also, there is an online distance between two points calculator at: Distance between two points Calculator http://easycalculation.com/analytical/distance.php

OpenStudy (anonymous):

nope...It isn't equilateral \[AB=\sqrt{17}\] \[BC=\sqrt{18}\] \[AC=\sqrt{17}\]

OpenStudy (anonymous):

AB=AC but not = BC

OpenStudy (anonymous):

as salmon said i dont think its a equilateral cuze d=squar root (x2-x1)^2+(y2-y1)^2 ab=(2,1) (3,5) squar root (3-2)^2+(5-1)^2= squar root 1+16= squar root 17 =4.123 bc=(3,5) (6,2) squar root (6-3)^2+(2-5)^2= squar root 9+9= squar root 18 = 4.242 ca=(6,2) (2,1) squar root (2-6)^2+(1-2)^2 squar root 16+1=squar root 17= 4.123

Directrix (directrix):

The next question is what do we do about this: If it is not an equilateral triangle, what changes can be made to make it equilateral? Be specific.

OpenStudy (anonymous):

how to do the 2nd question ??

OpenStudy (anonymous):

Actually their are number of ways to make it an equilateral...right??

OpenStudy (anonymous):

i dont know !

Directrix (directrix):

@Salmon >> Actually their are number of ways to make it an equilateral...right?? Such as -- Give us a specific example.

Directrix (directrix):

I don't know if we are supposed to use variations of the given coordinates to form an equilateral triangle. That is not stated as a requirement. If any three sets of coordinates pairs that form an equilateral triangle will suffice, then I think {(4,0); (-1, 5√3); and (-6,0)} are vertices of an equilateral triangle of side 10. @luka1998

OpenStudy (anonymous):

thx

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

I think you have change all the co-ordinates of the triangle...You should have only changed one of them to make it equilateral...